OpenStudy (anonymous):

Given the system of equations: y  =  3 x  +  9  4 x − 9 y  =  − 8 Find the  y-coordinate of the point of intersection of the two lines.

TheSmartOne (thesmartone):

do you know the method of substitution?

OpenStudy (anonymous):

method of substitution is when you plug one equation into another?

OpenStudy (emmatassone):

If you want to see the intersection point you have to look for the point that comply both equation. e.g. if I have this system: a) x+y=1 b) x-y=0 and i want to find now y-coordinate that comply the equations, so from b) we have x=y so we replace this information in a) then : x+x=1 ==> 2x=1 ====> x=1/2 since x=y ===> Y=1/2 in my example

TheSmartOne (thesmartone):

yes, you can plug in the first equation in to the second one

TheSmartOne (thesmartone):

and then solve for x

TheSmartOne (thesmartone):

then plug in the value you got for x into any of those equations and solve for y

OpenStudy (emmatassone):

you can work analogous in your problem

OpenStudy (anonymous):

which one is easier? I am not sure where emma got the 2x from here example. came from. If I plug one equation into another do I plug it into the front or back?

OpenStudy (emmatassone):

i used the substitution method xD i just plug the y=x in the other equation

OpenStudy (emmatassone):

you have to clear a variable first and then plug that variable in the other equation

OpenStudy (emmatassone):

I can do the example again exaplining it better if you want

OpenStudy (anonymous):

I am not sure where I put the y= 3x + 9 ... into 4x-9y=-8

OpenStudy (emmatassone):

*explaining

OpenStudy (emmatassone):

Here its another example with this system: \[4x+6y=0\] \[5x=3y+2\] So, what you have to do is clear one variable from any equation, im going to clear variable x from first equation:\[4x+6y=0 \rightarrow 4x=-6y \rightarrow x=\frac{ -6 }{ 4 }.y\] Once we cleared x, we plug it in the second equation:\[5x=3y+2\] replacing it: \[5(\frac{ -6 }{ 4 }.y)=3y+2\] \[\frac{ -30 }{ 4 }.y=3y+2\] \[-\frac{ 30 }{ 4 }y-3y=2\] \[-\frac{ 21 }{ 2 }y=2\] Finally:\[y=-\frac{ 4 }{ 21 }\]

OpenStudy (emmatassone):

Hope it help

OpenStudy (anonymous):

thank you